Zhdanov Inc. forecasts that its free cash flow in the coming year, that is, at t = 1, will be -$10 million, but its FCF at t = 2 will be $20 million. After Year 2, FCFis expected to grow at a constant rate of 4% forever. If the weighted average cost of capital is 14%, what is the firm's value of operations, in millions?
Answer #1
FCF1= -$10M, FCF2 = $20M
So here's what you need to do, before discounting the cash flows by the WACC (14%), calculate the terminal value, sometimes called the horizon value (thevalue after which the cash flows are constant), to do so, you need to calculate FCF3 for g=4% which we have FCF3 = FCF2*(1+g). ie FCF3= 20*(1+4%) =$20.8M

After doing so, calculate the terminal value using this formula: TV = FCF3/[WACC-g].
ie TV =20.8/(14%-4%) = $208M
Finally you just need to discount each FCF by the WACC, sum up the results (don't forget to add the TV to the last cash flow, so your last cash flow shouldbe FCF2/(1+WACC)+TV/(1+WACC)^2), after summing up the results, you got the firm's value.
= -10+20/(1+14%)^1 + 208/(1+14%)^2 = $ $167.59

So ans is $167M
answered by: connie